# Multimodal Needle in a Haystack: Benchmarking Long-Context Capability of Multimodal Large Language Models

### v1

@@ 1 Introduction
Recent breakthroughs in multimodal large language models (MLLMs) have enabled a wide range of applications, spanning from visual question answering to cross-modal retrieval [30, 28]. To evaluate the capabilities and limitations of MLLMs, various benchmarks have been proposed, focusing on challenges such as reasoning [30, 21, 20], perception [9, 29], and hallucination [11].
Despite significant progress, the evaluation of MLLMs for long-context understanding has been lagging. Current evaluation methods and benchmarks [30, 28, 19, 21, 9, 29, 6, 8, 20, 23] either (1) assume the use of single or limited images as inputs, failing to stress-test MLLMs’ long-context capabilities or (2) only contain a limited numbers of data points (referred to as “samples” in this paper), lacking in statistical significance and therefore often rendering the evaluation inconclusive. These gaps limit the development of MLLMs capable of effectively handling long-context hybrid-modality inputs, which is crucial for broader applications.
Figure 1: MMNeedle evaluation overview. Correct answers are marked with checkmark ($\checkmark$), while the incorrect answers are marked with cross ($\bm{\times}$). Our evaluation setup involves the following key components: (a) Needle Sub-Image: The needle sub-image to be retrieved based on the given caption. (b) Haystack Image Inputs: The long-context visual inputs consist of $M$ images, each stitched from $N\times N$ sub-images. (c) Text Inputs (Instructions and Caption): Detailed instructions to MLLMs, followed by a caption describing the needle, i.e., sub-image $20$. See Sec. 3.4 for MMNeedle’s complete instructions. (d) LLM Outputs: The answers from different MLLMs, indicating their ability to accurately locate the needle in the haystack based on the given caption. The expected output is composed of the model’s identification of the index, row, and column of the matching sub-image. The results showcase the comparative performance of various models: GPT-4o correctly predicts the exact location of the needle; Gemini Pro 1.5 only correctly predicts the image index of the needle; other API models predict incorrect locations; open-source models often output with wrong formats.
@@ 2 Related Work
Existing benchmarks for MLLMs mainly focus on limited image inputs, such as reasoning [30, 21, 20], perception [9, 29], hallucination [11], where the answers are based on either single or only a handful of images. They are therefore not suitable for evaluating MLLMs’ long-context capability for visual inputs. Recent work [10, 13, 15, 31] on LLMs employs the needle-in-a-haystack test [12] to evaluate the long-context capability of large language models (LLMs), where the LLM is expected to answer the question by finding the corresponding information among a long irrelevant corpus as context. However, these datasets and benchmarks are not applicable for the multimodal setting. Google’s technical report [23] has showcased Gemini 1.5 Pro’s capability of finding the needle in an audio or video haystack. However, its evaluation (1) involves only one single sample rather than a complete dataset, obviously lacking statistical significance and therefore rendering the evaluation inconclusive [^1] Our MMNeedle results show that Gemini 1.5 Pro’s performance does drop a lot with long contexts, especially with multiple sub-images in the same image. This demonstrates the value of our MMNeedle as a comprehensive benchmark., and (2) does not involve a large set of unrelated images, which is the focus of MMNeedle. There is also work on the retrieval capability of small objects in a single large image [22] or retrieval from large external image datasets [5], but none of them are concerned with in-context image retrieval, particularly for long-context multimodal evaluation.
In contrast to existing benchmarks, our MMNeedle benchmark includes a dataset of $40{,}000$ images, $560{,}000$ captions, and $280{,}000$ needle-haystack pairs (more details in Sec. 3), rather than only one (or a handful of) needle-haystack pair(s) [12, 23]. MMNeedle also includes a diverse set of metrics and evaluation protocols, covering different numbers of needle sub-images and needle sub-images. These differences set MMNeedle apart from existing benchmarks and are essential to evaluate MLLMs’ long-context capability comprehensively.
@@ 3.1 Overview
Evaluation Goals. As illustrated in Fig. 1, our MMNeedle aims to evaluate the MLLMs’ three key capabilities within one forward pass: (1) understanding the semantics of both visual and textual inputs, (2) retrieving the sub-image (needle) from long-context images (haystack), and (3) understanding and following the instructions [26] to output the location of the sub-image (needle) in the correct format.
@@ 3.2 MMNeedle Dataset
  • More Images: We increase the number of images in the inputs for MLLMs to extend the visual context length. Specifically, we use two different numbers of images $M$ in the prompt, i.e., $M=1$ or $M=10$. Note that we choose $M=10$ because it is the largest number of input images that GPT-4V/GPT-4o and LLaVA-Llama-3 can support.
  • Image Stitching: We stitch small images into a single large image as the input to MLLMs. Specifically, we use $N\times N$ sub-images ($N\in\{1,2,4,8\}$) to compose a stitched image with $N$ rows and $N$ columns, each combination of row and column indices (r,c) corresponding to a sub-image. Fig. 1(b) shows an example of $2\times 2$ stitching, with $4$ sub-images in $1$ stitched image.
Purpose of Image Stitching. The purpose of image stitching is to: (1) Extend the effective context length. For example, stitching $M=10$ images, each with $N\times N=8\times 8$ sub-images, results in a long context of $640$ sub-images. This setup tests MLLMs’ long-context capabilities. (2) Test MLLMs’ localization capability by requiring them to pinpoint sub-images within a large image based on specific captions. Combining both dimensions provides comprehensive settings for our evaluation: $(M,N)=(1,2),(1,4),(1,8),(10,1),(10,2),(10,4),(10,8)$. Note that $(M,N)=(1,1)$ is excluded, as finding an image within a single image is trivial.
@@
Image Data. In this paper, we use the MS COCO 2014 validation set [17] as our source dataset for constructing our MMNeedle dataset. Note that our data construction approach is agnostic to the dataset and can be applied to any dataset containing images with paired captions that describe the content of the images. We resize each original image from the MS COCO 2014 validation set to $256\times 256$ pixels before stitching them into a larger image. The image resolution of $256$ pixels is chosen to ensure sufficient image quality; our preliminary studies show that humans (and MLLMs) cannot effectively recognize MS COCO images with resolution lower than $256$ (see examples in Fig. 1 and more in the Supplement). We then stitch these sub-images using stitching sizes of $1\times 1$, $2\times 2$, $4\times 4$, and $8\times 8$, leading to larger images with resolutions of $256\times 256$, $512\times 512$, $1024\times 1024$, and $2048\times 2048$, respectively. Given that Claude 3 supports a maximum resolution of $1092\times 1092$ pixels and GPT-4 (including GPT-4V and GPT-4o) supports a maximum resolution of 2000 pixels for the long side of an image, we have chosen 2048 pixels as the maximum resolution for our stitched images. Note that these models will resize images that exceed their respective size limits.
@@ 3.3 Dataset Construction: Automated Sampling
3.4 Prompt Design
For single-needle evaluation, we use the following prompt for the evaluated LLM:
[…]
where the instructions to MLLMs is as follows:
Given M images indexed from 1 to M, each divided into N*N sub-images,
identify the sub-image that best matches the provided caption.
Respond with ‘‘index, row, column’’ and nothing else. For example,
‘‘1, 2, 3’’ indicates the sub-image in the first image, second row,
and third column. If no match is found, respond only with ‘‘-1’’.
We use a similar prompt for the multi-needle setting (see the Supplement for details).
@@ 3.5 Evaluation Metrics
  • Exact Accuracy (success rate of the needle retrieval [23]) is the proportion of samples where the model correctly predicts the needle sub-image’s location, i.e., index $m$, row $r$ and column $c$.
Multiple Needles. We use similar metrics for the multi-needle setting (details in the Supplement).
@@
  • Ground Truth Format. (1) For each positive sample, i.e., the needle sub-image is in the context, the ground-truth output is “$m,r,c$” that describes the location of the needle, where $m$ is the image index ($m\in{1,...,M}$), and $r,c$ are the row and column of the sub-image (needle) in image $m$, respectively ($r,c\in{1,...,N}$). (2) For each negative sample, i.e., no needle sub-image is in the context, the ground-truth output is “-1”, indicating the needle does not exist. The multi-needle setting uses a similar format (details in the Supplement).
@@ 4.1 Evaluated MLLMs
  • API-Based Models. We evaluate state-of-the-art API-based MLLMs, including Claude 3 Opus (Feb 2024) [1], Gemini Pro 1.0 (Feb 2024) [24], Gemini Pro 1.5 (May 2024) [23], GPT-4V (March 2024) [3], and GPT-4o (May 2024) [2].
  • Open-Source Models. We evaluate top open-source multimodal LLMs, including CogVLM (CogVLM-17B/CogVLM2-Llama-3) [25], Fuyu-8B [4], mPLUG-Owl-v2 [27], InstructBLIP (InstructBLIP-Vicuna-13B/InstructBLIP-Flan-T5-XXL) [7], IDEFICS2 [14], and LLaVA-Llama-3 [16]. Note that CogVLM and InstructBLIP do not support multi-image inputs; therefore, we do not test them for our multi-image ($M=10$) settings.
See the Supplement for more details on evaluated MLLMs.
@@ 4.2 Overview of MMNeedle Evaluation Results
See Sec. 4.3 below for details and more evaluation results.
@@
API-Based Models
Claude 3 Opus 75.38 74.77 52.25 58.70 58.00 12.30 56.36 54.85 1.60
@@
Open-Source Models
CogVLM-17B 99.90 0.80 0.00 97.50 3.30 0.10 96.90 22.90 0.30
@@
4.3 Detailed Results of the Three Defined Metrics
In this section, we discuss the results of the MMNeedle evaluation in various settings of $(M,N,K)$ across three metrics: Existence, Index, and Exact Accuracy, as defined in Sec. 3.5. More results are available in the Supplement.
@@ 4.3 Detailed Results of the Three Defined Metrics
API-Based Models
Claude 3 Opus 83.77 67.23 66.93 66.60 9.90 4.60 64.78 6.46 0.40 54.13 5.93 0.00
@@
Open-Source Models
@@
Stitching $2\times 2$ $4\times 4$ $8\times 8$
API-Based Models
Claude 3 Opus 100.00 66.00 32.00 97.00 31.00 1.00 98.00 25.00 0.00
Gemini Pro 1.0 100.00 79.80 9.09 95.00 50.00 2.00 68.00 11.00 0.00
Gemini Pro 1.5 100.00 94.95 87.88 100.00 84.00 22.00 98.00 80.00 6.00
GPT-4V 100.00 90.72 71.13 100.00 95.00 34.00 100.00 93.41 1.10
GPT-4o 100.00 84.00 76.00 100.00 84.00 57.00 100.00 78.00 2.00
Open-Source Models
CogVLM-17B 100.00 0.00 0.00 100.00 0.00 0.00 100.00 0.00 0.00
CogVLM2-Llama-3 100.00 0.00 0.00 100.00 0.00 0.00 100.00 0.00 0.00
Fuyu-8B 100.00 0.00 0.00 100.00 0.00 0.00 100.00 0.00 0.00
mPLUG-Owl-v2 98.00 0.00 0.00 94.00 2.00 0.00 96.00 3.00 0.00
IDEFICS2-8B 100.00 0.00 0.00 100.00 0.00 0.00 100.00 0.00 0.00
LLaVA-Llama-3 100.00 0.00 0.00 100.00 2.00 0.00 100.00 12.00 0.00
@@
Results on Multi-Needle Samples ($K>1$). Table 4 shows the results of different models on multi-needle samples, i.e., the number of needles $K=2$. Gemini Pro 1.5 achieves the highest exact accuracy $87.88\%$ on $2\times 2$ samples, and GPT-4o achieves the highest exact accuracy $57.00\%$ on $4\times 4$ samples. In contrast, the exact accuracy of open-source models is close to zero for all stitching sizes. These results indicate a large gap between the API-based and the open-source models. See the Supplement for more results and analysis on multi-needle samples ($K=2$ or $K=5$).

### v2

@@ Abstract
Multimodal Needle in a Haystack: Benchmarking Long-Context Capability of Multimodal Large Language Models
Hengyi Wang1thanks: Correspondence to: Hengyi Wang <hengyi.wang@rutgers.edu> ,  Haizhou Shi1,  Shiwei Tan1,  Weiyi Qin1,  Wenyuan Wang1, Tunyu Zhang1,  Akshay Nambi2,  Tanuja Ganu2,  Hao Wang1 1Rutgers University, 2Microsoft Research https://mmneedle.github.io
@@ 1 Introduction
Recent breakthroughs in multimodal large language models (MLLMs) have enabled a wide range of applications, spanning from visual question answering to cross-modal retrieval Yue et al. (2023); Ying et al. (2024). To evaluate the capabilities and limitations of MLLMs, various benchmarks have been proposed, focusing on challenges such as reasoning Yue et al. (2023); Padlewski et al. (2024); Lu et al. (2023), perception Fu et al. (2024b); Yu et al. (2023), and hallucination Guan et al. (2023).
Despite significant progress, the evaluation of MLLMs for long-context understanding has been lagging. Current evaluation methods and benchmarks Yue et al. (2023); Ying et al. (2024); Liu et al. (2023); Padlewski et al. (2024); Fu et al. (2024b); Yu et al. (2023); Chen et al. (2024); Fu et al. (2024a); Lu et al. (2023); Reid et al. (2024) either (1) assume the use of single or limited images as inputs, failing to stress-test MLLMs’ long-context capabilities or (2) only contain a limited numbers of data points (referred to as “samples” in this paper), lacking in statistical significance and therefore often rendering the evaluation inconclusive. These gaps limit the development of MLLMs capable of effectively handling long-context hybrid-modality inputs, which is crucial for broader applications.
Figure 1: MMNeedle evaluation overview. Correct answers are marked with checkmark ($\checkmark$), while the incorrect answers are marked with cross ($\bm{\times}$). Our evaluation setup involves the following key components: (a) Needle Sub-Image: The needle sub-image to be retrieved based on the given caption. (b) Haystack Image Inputs: The long-context visual inputs consist of $M$ images, each stitched from $N\times N$ sub-images. (c) Text Inputs (Instructions and Caption): Detailed instructions to MLLMs, followed by a caption describing the needle, i.e., sub-image $20$. See Sec. A for MMNeedle’s complete instructions. (d) LLM Outputs: The answers from different MLLMs, indicating their ability to accurately locate the needle in the haystack based on the given caption. The expected output is composed of the model’s identification of the index, row, and column of the matching sub-image. The results showcase the comparative performance of various models: GPT-4o correctly predicts the exact location of the needle; Gemini Pro 1.5 only correctly predicts the image index of the needle; other API models predict incorrect locations; open-source models often output with wrong formats.
@@ 2 Related Work
Existing benchmarks for MLLMs mainly focus on limited image inputs, such as reasoning Yue et al. (2023); Padlewski et al. (2024); Lu et al. (2023); Song et al. (2024), perception Fu et al. (2024b); Yu et al. (2023), hallucination Guan et al. (2023), where the answers are based on either single or only a handful of images. They are therefore not suitable for evaluating MLLMs’ long-context capability for visual inputs. Recent work Fu et al. (2024c); Kuratov et al. (2024); Levy et al. (2024); Zhao et al. (2024) on LLMs employs the needle-in-a-haystack test Kamradt (2023) to evaluate the long-context capability of large language models (LLMs), where the LLM is expected to answer the question by finding the corresponding information among a long irrelevant corpus as context. However, these datasets and benchmarks are not applicable for the multimodal setting. Google’s technical report Reid et al. (2024) has showcased Gemini 1.5 Pro’s capability of finding the needle in an audio or video haystack. However, its evaluation (1) involves only one single sample rather than a complete dataset, obviously lacking statistical significance and therefore rendering the evaluation inconclusive  [^1] Our MMNeedle results show that Gemini 1.5 Pro’s performance does drop a lot with long contexts, especially with multiple sub-images in the same image., and (2) does not involve a large set of unrelated images, which is the focus of MMNeedle. There is also work on the retrieval capability of small objects in a single large image Pawlowski et al. (2019) or retrieval from large external image datasets Brogan et al. (2019), but none of them are concerned with in-context image retrieval, particularly for long-context multimodal evaluation.
In contrast to existing benchmarks, our MMNeedle benchmark includes a dataset of $40{,}000$ images, $560{,}000$ captions, and $280{,}000$ needle-haystack pairs (more details in Sec. 3), rather than only one (or a handful of) needle-haystack pair(s) Kamradt (2023); Reid et al. (2024). MMNeedle also includes a diverse set of metrics and evaluation protocols, covering different numbers of needle sub-images and needle sub-images. These differences set MMNeedle apart from existing benchmarks and are essential to evaluate MLLMs’ long-context capability comprehensively.
@@ 3.1 Overview
Evaluation Goals. As illustrated in Fig. 1, our MMNeedle aims to evaluate the MLLMs’ three key capabilities within one forward pass: (1) understanding the semantics of both visual and textual inputs, (2) retrieving the sub-image (needle) from long-context images (haystack), and (3) understanding and following the instructions Xia et al. (2024) to output the location of the sub-image (needle) in the correct format.
@@ 3.2 MMNeedle Dataset
  • More Images: We increase the number of images in the inputs for MLLMs to extend the visual context length. Specifically, we use two different numbers of images $M$ in the prompt, i.e., $M=1$ or $M=10$. Note that we choose $M=10$ because it is the largest number of input images that GPT-4V/GPT-4o can support (see Table 1 and Appendix A).
  • Image Stitching: We stitch small images into a single large image as the input. Specifically, we use $N\times N$ sub-images ($N\in\{1,2,4,8\}$) to compose a stitched image with $N$ rows and $N$ columns, each combination of row and column indices (r,c) corresponding to a sub-image. Fig. 1(b) shows an example of $2\times 2$ stitching, with $4$ sub-images in $1$ stitched image.
Table 1: Maximum numbers of images per request for Azure GPT-4V/o , OpenAI GPT-4V/o, Claude, and Gemini. "*" indicates that the OpenAI GPT-4V/o API supports at most $10$ images with high quality. Other numbers are hard limits. See Appendix A for details.
Model GPT-4 (Az.) GPT-4 (Op.) Claude Gemini
Limit 10 10 20 16
Purpose of Image Stitching. The purpose of image stitching is to: (1) Extend the effective context length. For example, stitching $M=10$ images, each with $N\times N=8\times 8$ sub-images, results in a long context of $640$ sub-images. This setup tests MLLMs’ long-context capabilities. (2) Test MLLMs’ localization capability by requiring them to pinpoint sub-images within a large image based on specific captions. For details, see Appendix A.
Combining both dimensions provides comprehensive settings for our evaluation: $(M,N)=(1,2),(1,4),(1,8),(10,1),(10,2),(10,4),(10,8)$. Note that $(M,N)=(1,1)$ is excluded, as finding an image within a single image is trivial. Note that MMNeedle covers typical, real-world MLLM use-cases. Specifically, single, complete images correspond to our setting with the number of images $M=10$ and the stitch size $N\times N=1\times 1$.
@@
Image Data. In this paper, we use the MS COCO 2014 validation set Lin et al. (2014) as our source dataset for constructing our MMNeedle dataset. Note that our data construction approach is agnostic to the dataset and can be applied to any dataset containing images with paired captions that describe the content of the images. We resize each original image from the MS COCO 2014 validation set to $256\times 256$ pixels before stitching them into a larger image. The image resolution of $256$ pixels is chosen to ensure sufficient image quality; our preliminary studies show that humans (and MLLMs) cannot effectively recognize MS COCO images with resolution lower than $256$ (see examples in Fig. 1 and more in Appendix A). We then stitch these sub-images using stitching sizes of $1\times 1$, $2\times 2$, $4\times 4$, and $8\times 8$, leading to larger images with resolutions of $256\times 256$, $512\times 512$, $1024\times 1024$, and $2048\times 2048$, respectively. Given that Claude 3 supports a maximum resolution of $1092\times 1092$ pixels and GPT-4 (including GPT-4V and GPT-4o) supports a maximum resolution of 2000 pixels for the long side of an image, we have chosen 2048 pixels as the maximum resolution for our stitched images. Note that these models will resize images that exceed their respective size limits.
@@ 3.4 Evaluation Metrics
  • Exact Accuracy (success rate of the needle retrieval Reid et al. (2024)) is the proportion of samples where the model correctly predicts the needle sub-image’s location, i.e., index $m$, row $r$ and column $c$.
Multiple Needles. We use similar metrics for the multi-needle setting (details in Appendix B).
@@
  • Ground Truth Format. (1) For each positive sample, i.e., the needle sub-image is in the context, the ground-truth output is “$m,r,c$” that describes the location of the needle, where $m$ is the image index ($m\in{1,...,M}$), and $r,c$ are the row and column of the sub-image (needle) in image $m$, respectively ($r,c\in{1,...,N}$). (2) For each negative sample, i.e., no needle sub-image is in the context, the ground-truth output is “-1”, indicating the needle does not exist. The multi-needle setting uses a similar format (details in Appendix B).
@@ 4.1 Evaluated MLLMs
  • API-Based Models. We evaluate state-of-the-art API-based MLLMs, including Claude 3 Opus (Feb 2024) ant (2023), Gemini Pro 1.0 (Feb 2024) Team et al. (2023), Gemini Pro 1.5 (May 2024) Reid et al. (2024), GPT-4V (March 2024) Achiam et al. (2023), and GPT-4o (May 2024) ope (2024).
  • Open-Source Models. We evaluate top open-source multimodal LLMs, including CogVLM (CogVLM-17B/CogVLM2-Llama-3) Wang et al. (2023), Fuyu-8B Bavishi et al. (2023), mPLUG-Owl-v2 Ye et al. (2023), InstructBLIP (InstructBLIP-Vicuna-13B/InstructBLIP-Flan-T5-XXL) Dai et al. (2024), IDEFICS2 Laurençon et al. (2024), and LLaVA-Llama-3 Li et al. (2024). Note that CogVLM and InstructBLIP do not support multi-image inputs; therefore, we do not test them for our multi-image ($M=10$) settings.
See Appendix C for more details on evaluated MLLMs.
@@ 4.2 Overview of MMNeedle Evaluation Results
We also analyze the error patterns. As illustrated in Fig. 2, the models demonstrate higher accuracy when the needles are positioned in the corners of the image compared to when they are located in the center. This trend is particularly pronounced in Gemini-1.5 and LLaVA-Llama-3, in contrast to GPT-4o. See Sec. 4.3 below for details and more evaluation results.
@@
API-Based Models Claude 3 Opus 75.38 74.77 52.25 58.70 58.00 12.30 56.36 54.85 1.60
@@
Open-Source Models CogVLM-17B 99.90 0.80 0.00 97.50 3.30 0.10 96.90 22.90 0.30
@@
API-Based Models Claude 3 Opus 83.77 67.23 66.93 66.60 9.90 4.60 64.78 6.46 0.40 54.13 5.93 0.00
@@
Open-Source Models Fuyu-8B 100.00 0.00 0.00 100.00 0.00 0.00 100.00 0.00 0.00 100.00 0.00 0.00
@@
4.3 Detailed Results of the Three Defined Metrics
In this section, we discuss the results of the MMNeedle evaluation in various settings of $(M,N,K)$ across three metrics: Existence, Index, and Exact Accuracy, as defined in Sec. 3.4. More results are available in Appendix D.
@@ 4.3 Detailed Results of the Three Defined Metrics
Stitching $2\times 2$ $4\times 4$ $8\times 8$
API-Based Models Claude 3 Opus 100.00 66.00 32.00 97.00 31.00 1.00 98.00 25.00 0.00
Gemini Pro 1.0 100.00 79.80 9.09 95.00 50.00 2.00 68.00 11.00 0.00
Gemini Pro 1.5 100.00 94.95 87.88 100.00 84.00 22.00 98.00 80.00 6.00
GPT-4V 100.00 90.72 71.13 100.00 95.00 34.00 100.00 93.41 1.10
GPT-4o 100.00 84.00 76.00 100.00 84.00 57.00 100.00 78.00 2.00
Open-Source Models CogVLM-17B 100.00 0.00 0.00 100.00 0.00 0.00 100.00 0.00 0.00
CogVLM2-Llama-3 100.00 0.00 0.00 100.00 0.00 0.00 100.00 0.00 0.00
Fuyu-8B 100.00 0.00 0.00 100.00 0.00 0.00 100.00 0.00 0.00
mPLUG-Owl-v2 98.00 0.00 0.00 94.00 2.00 0.00 96.00 3.00 0.00
IDEFICS2-8B 100.00 0.00 0.00 100.00 0.00 0.00 100.00 0.00 0.00
LLaVA-Llama-3 100.00 0.00 0.00 100.00 2.00 0.00 100.00 12.00 0.00
@@
Results on Multi-Needle Samples ($K>1$). Table 4 shows the results of different models on multi-needle samples, i.e., the number of needles $K=2$. Gemini Pro 1.5 achieves the highest exact accuracy $87.88\%$ on $2\times 2$ samples, and GPT-4o achieves the highest exact accuracy $57.00\%$ on $4\times 4$ samples. In contrast, the exact accuracy of open-source models is close to zero for all stitching sizes. These results indicate a large gap between the API-based and the open-source models. See Appendix D for more results and analysis on multi-needle samples ($K=2$ or $K=5$).
@@
Summary. These results show that our existence, index, and exact accuracy are designed to differentiate the model capabilities across various settings while also facilitating a transition from easier to more challenging tasks.
For example, we demonstrate that various metrics highlight the long-context capabilities of models under different settings:
  • Exact Accuracy: In Table 2, where the number of input images $M=1$, we focus on evaluating exact accuracy, which measures whether the model correctly predicts both the row and column of the needle.
  • Index Accuracy: In Table 3, where the number of input images $M=10$, we emphasize index accuracy, assessing whether the model correctly identifies the image index within the image haystack. Together with Exact Accuracy, it is crucial for evaluating whether an MLLM can understand images and sub-images in the long-context scenario.
  • Existence Accuracy: In Table 4, where negative samples are introduced, we evaluate existence accuracy, which reflects whether the model correctly determines that the needle is not present in the haystack. This is particularly relevant for benchmarking hallucination in MLLMs.
These analyses underscore the different use cases and the necessity of our coarse-to-fine metrics.